# Almost Sure Convergence for the Last Iterate of Stochastic Gradient Descent Schemes

### v2

@@ 4.2. Error Rate with High Probability
If all assumptions in proposition 4.5 are satisfied, then under the event $\pazocal{E}_{t}^{\delta}$, with the notation in proposition 4.5 and lemma B.6,
@@
Under the assumptions of corollary 4.6, $T>0$ is fixed and $t\leq T$. Therefore, $\pazocal{E}_{k}$ is in $\mathscr{F}_{t}$. The proof follows directly from proposition 4.5, lemma B.6.2, B.6.3b, and B.6.3c. ∎
@@
where $\xi_{t}$ and $\varphi_{t}$ are defined in (4.11). By lemma B.6.3b, we have
@@
It remains to estimate the martingale part in (4.21). Applying Cauchy-Schwarz and lemma B.6.3b gives
@@
By Cauchy-Schwarz, $\gamma$-Hölder property of $\nabla F$, lemma B.6.3b, for all $t\in[\tau,T]$,
@@
We apply Cauchy-Schwarz and lemma B.6.3b to get
@@
Case $\beta\in(0,1)$. The recursion of $\|v_{t}\|^{2}$ in the previous paragraph can be rewritten into an integral equation by proposition C.1:
@@
We estimate the parentheses in the first term. By lemma B.6.3c and the convexity of $F$, we get
@@
Now, examine the second term on the right hand side of (4.27). By lemma B.6.3a, we have $\alpha_{t}^{2}\,\|\nabla\ell(Z_{t},w_{t})\|^{2}\leq\alpha_{t}^{2}\,(a_{1}\ell(Z_{t},w_{t})+a_{2})$. By lemma 4.1, we get $\sum\alpha_{t}^{2}\,\ell(Z_{t},w_{t})<K_{2}$. Therefore, using the expression for $X_{t}$ in (4.28), under the event $\pazocal{E}_{T}^{\prime}$,
@@
Also, since $v_{1}=0$, we have $\langle\nabla F(w_{1}),\,v_{1}\rangle=0$. Using proposition C.1 and switching the order of integration, under the event $\pazocal{E}$, the following occurs with probability $\mathbb{P}(\pazocal{E})\geq 1-\delta$ for all $\tau>1$:
@@
Since $Z_{t}\in\ell^{1}(\mathbb{N})$, the above is a Gronwall-type inequality. By proposition C.2, we get
@@
We require that r1<pr-1<-p and $p(1+\gamma)+r<\min(-p,-1)$ so that the second and third term on the right hand side are both summable. Since $p\in(\frac{1}{1+\gamma},1)$, such condition holds if and only if $r<\min(1-p,p(1+\gamma)-1)$. Now, applying proposition 2.4 yields $\limsup_{t\rightarrow\infty}\,\mathbb{E}[t^{r}\cdot Y_{t}\,|\,\mathscr{F}_{s}]<\infty$.
@@
$\mathbb{E}[T^{r}\cdot Y_{T}\,|\,\mathscr{F}_{s}]-s^{r}\cdot Y_{s}$ $\lesssim\mathbb{E}\left[\sum_{t=s}^{T-1}(\alpha_{t}^{1+\gamma}\cdot t^{r}+t^{r-1})\cdot Y_{t}\bigg|\mathscr{F}_{s}\right]+\mathbb{E}\left[\sum_{t=s}^{T-1}t^{-p(1+\gamma)+r}Z_{t}\bigg|\mathscr{F}_{s}\right]$ (A.3)
$\leq\mathbb{E}\left[\sum_{t=1}^{\infty}(\alpha_{t}^{1+\gamma}\cdot t^{r}+t^{r-1})\cdot Y_{t}\bigg|\mathscr{F}_{s}\right]+\mathbb{E}\left[\sum_{t=1}^{\infty}t^{-p(1+\gamma)+r}Z_{t}\bigg|\mathscr{F}_{s}\right].$ (A.4)
@@
This implies that $\lim_{s\uparrow\infty}s^{r}\cdot Y_{s}$ is $\mathscr{F}_{\infty}$-measurable. Furthermore, since $\sum_{t=1}^{\infty}\alpha_{t}\,Y_{t}<\infty$ almost surely, by [7, theorem 3.3.1], we have $\min_{s\leq t}Y_{s}=o(t^{p-1})$. Therefore, taking $\liminf_{s\uparrow\infty}$ on both sides of (A.4) yields almost surely
@@
Note that applying the expectation to (A.2) and proposition 2.4 yields $\sup_{t\geq 0}\,\mathbb{E}[t^{r}\cdot Y_{t}]<\infty$. If we take $0<s<T$ and $s\uparrow\infty$ so that $T\uparrow\infty$ and use DCT on $\lim_{T\uparrow\infty}\mathbb{E}[T^{r}Y_{T}\,|\,\mathscr{F}_{s}]$ and the right hand side of (A.3), we get
@@
Appendix B Some Facts Used in Proofs
@@
Proof of (1). We use the same argument as (B.1), but replace $\ell(z,w_{*})$ with $F_{*}$.
Proof of (2). Use $F(w_{t})=F(w_{t})-F_{*}+F_{*}$ and apply lemma B.4.
Proof of (3.a). Applying lemma B.4 and subsequently Young’s inequality [8, eq 4.3] with $p=(1+\gamma)/(1-\gamma)$ and $q=(1+\gamma)/(2\gamma)$ for $\gamma\in(0,1]$ yields
$\|\nabla\ell(z,w)\|^{2}$ $\leq\left(\frac{1+\gamma}{\gamma}\right)^{\frac{2\gamma}{1+\gamma}}L^{\frac{2}{1+\gamma}}\left(\ell(z,w)-\ell(z,w_{*})\right)^{\frac{2\gamma}{1+\gamma}}$ (B.1)
@@
Appendix C Gronwall-type Inequalities
Proposition C.1.
is the unique solution to the difference equation
$\begin{cases}Y_{n_{0}+n}=\beta^{2}\,Y_{n_{0}+n-1}+X_{n_{0}+n-1}~,~n>0.\\ Y_{n_{0}}=y.\end{cases}$ (C.1)
Proof.
$=X_{n_{0}+n}+\beta^{2(n+1)}y+\sum_{k=1}^{n}\beta^{2k}X_{n_{0}+n-k}$
$=Y_{n_{0}+n+1}.$
For uniqueness, if $Z_{n_{0}+n}$ satisfies (C.1), then $Z_{n_{0}+n}-Y_{n_{0}+n}=\beta^{2}(Z_{n_{0}+n-1}-Y_{n_{0}+n-1})$ for all $n>0$ with $Z_{n_{0}}-Y_{n_{0}}=0$. This implies $Z_{n_{0}+n}-Y_{n_{0}+n}=(Z_{n_{0}}-Y_{n_{0}})\,\beta^{2n}=0$ for all $n\geq 0$.
We can apply the previous claim and get $Z_{n_{0}+n}\leq-Y_{n_{0}+n}$. ∎
Proposition C.2.
Let $A\geq 0$, $X_{n}\geq 0$, $c_{n}\geq 0$.
Proof.
Define the sequence
$Y_{n}=A+\sum_{k=1}^{n}c_{k-1}\,X_{k-1}~~~,~~~Y_{0}=A$
$\frac{Y_{n}}{Y_{n-1}}\leq 1+c_{n-1}~~,~~\forall n\geq 1.$

### v3

@@ 4.2. Error Rate with High Probability
If all assumptions in proposition 4.5 are satisfied, then under the event $\pazocal{E}_{t}^{\delta}$, with the notation in proposition 4.5 and lemma C.6,
@@
Under the assumptions of corollary 4.6, $T>0$ is fixed and $t\leq T$. Therefore, $\pazocal{E}_{k}$ is in $\mathscr{F}_{t}$. The proof follows directly from proposition 4.5, lemma C.6.2, C.6.3b, and C.6.3c. ∎
@@
where $\xi_{t}$ and $\varphi_{t}$ are defined in (4.11). By lemma C.6.3b, we have
@@
It remains to estimate the martingale part in (4.21). Applying Cauchy-Schwarz and lemma C.6.3b gives
@@
By Cauchy-Schwarz, $\gamma$-Hölder property of $\nabla F$, lemma C.6.3b, for all $t\in[\tau,T]$,
@@
We apply Cauchy-Schwarz and lemma C.6.3b to get
@@
Case $\beta\in(0,1)$. The recursion of $\|v_{t}\|^{2}$ in the previous paragraph can be rewritten into an integral equation by proposition B.1:
@@
We estimate the parentheses in the first term. By lemma C.6.3c and the convexity of $F$, we get
@@
Now, examine the second term on the right hand side of (4.27). By lemma C.6.3a, we have $\alpha_{t}^{2}\,\|\nabla\ell(Z_{t},w_{t})\|^{2}\leq\alpha_{t}^{2}\,(a_{1}\ell(Z_{t},w_{t})+a_{2})$. By lemma 4.1, we get $\sum\alpha_{t}^{2}\,\ell(Z_{t},w_{t})<K_{2}$. Therefore, using the expression for $X_{t}$ in (4.28), under the event $\pazocal{E}_{T}^{\prime}$,
@@
Also, since $v_{1}=0$, we have $\langle\nabla F(w_{1}),\,v_{1}\rangle=0$. Using proposition B.1 and switching the order of integration, under the event $\pazocal{E}$, the following occurs with probability $\mathbb{P}(\pazocal{E})\geq 1-\delta$ for all $\tau>1$:
@@
Since $Z_{t}\in\ell^{1}(\mathbb{N})$, the above is a Gronwall-type inequality. By proposition B.2, we get
@@
We require that r1<pr-1<-p and $p(1+\gamma)+r<\min(-p,-1)$ so that the second and third term on the right hand side are both summable. Since $p\in(\frac{1}{1+\gamma},1)$, such condition holds if and only if $r<\min(1-p,p(1+\gamma)-1)$. Now, applying proposition 2.4 yields $\limsup_{t\rightarrow\infty}\,\mathbb{E}[t^{r}\cdot Y_{t}\,|\,\mathscr{F}_{s}]<\infty$. Note that under the same condition on $r$, we may apply expectation on both sides of (A.2) and then apply proposition 2.4 to conclude $\sup_{t\geq 0}\,\mathbb{E}[t^{r}\cdot Y_{t}]<\infty$.
@@
$\mathbb{E}[T^{r}\cdot Y_{T}\,|\,\mathscr{F}_{s}]-s^{r}\cdot Y_{s}$ $\lesssim\mathbb{E}\left[\sum_{t=s}^{T-1}(\alpha_{t}^{1+\gamma}\cdot t^{r}+t^{r-1})\cdot Y_{t}\bigg|\mathscr{F}_{s}\right]+\mathbb{E}\left[\sum_{t=s}^{T-1}t^{-p(1+\gamma)+r}Z_{t}\bigg|\mathscr{F}_{s}\right].$ (A.3)
$\mathbb{E}[T^{r}\cdot Y_{T}\,|\,\mathscr{F}_{s}]-s^{r}\cdot Y_{s}$ $\leq\mathbb{E}\left[\sum_{t=1}^{\infty}(\alpha_{t}^{1+\gamma}\cdot t^{r}+t^{r-1})\cdot Y_{t}\bigg|\mathscr{F}_{s}\right]+\mathbb{E}\left[\sum_{t=1}^{\infty}t^{-p(1+\gamma)+r}Z_{t}\bigg|\mathscr{F}_{s}\right].$ (A.4)
@@
This implies that $\lim_{s\uparrow\infty}s^{r}\,Y_{s}$ is $\mathscr{F}_{\infty}$-measurable. Furthermore, since $\sum_{t=1}^{\infty}\alpha_{t}\,Y_{t}<\infty$ almost surely, by [7, theorem 3.3.1], we have $\min_{s\leq t}Y_{s}=o(t^{p-1})$.
We need to conclude $\liminf_{s\uparrow\infty}s^{r}Y_{s}=0$ a.s.-$\mathbb{P}$. This holds if we require $Y_{t}>0$ for all $t\geq 0$ (counterexample can be constructed if otherwise) assuming $t^{r}\min_{s\leq t}Y_{s}\downarrow 0$ as $t\uparrow\infty$. If $Y_{t}>0$ for all $t\geq 0$, then one can take a subsequence $Y_{t_{n}}\downarrow 0$ decreasing to zero. For instance, one can take $\{t_{n}:n\in\mathbb{N}\}$ such that $Y_{t_{n}}=\min_{t\leq t_{n}}Y_{t}$. Therefore, we have $t_{n}^{r}Y_{t_{n}}=t_{n}^{r}\min_{t\leq t_{n}}Y_{t}$ where the RHS converges to zero. This implies that there exists a subsequence of $t^{r}Y_{t}$ converging to zero, i.e. $\liminf_{s\uparrow\infty}s^{r}Y_{s}=0$ a.s.-$\mathbb{P}$. Therefore, taking $\liminf_{s\uparrow\infty}$ on both sides of (A.4) yields almost surely
@@
Since $\mathbb{E}[\lim_{t\uparrow\infty}t^{r}Y_{t}\,|\,\mathscr{F}_{s}]$ converges as $s\uparrow\infty$, we know
[…]
If we take $0<s<T$ and $s\uparrow\infty$ so that $T\uparrow\infty$ and use DCT on $\lim_{T\uparrow\infty}\mathbb{E}[T^{r}Y_{T}\,|\,\mathscr{F}_{s}]$ and the right hand side of (A.3), we get
@@
Appendix B Gronwall-type Inequalities
@@
is the unique solution to the difference equation
$\begin{cases}Y_{n_{0}+n}=\beta^{2}\,Y_{n_{0}+n-1}+X_{n_{0}+n-1}~,~n>0.\\ Y_{n_{0}}=y.\end{cases}$ (B.1)
Proof.
$=X_{n_{0}+n}+\beta^{2(n+1)}y+\sum_{k=1}^{n}\beta^{2k}X_{n_{0}+n-k}$
$=Y_{n_{0}+n+1}.$
For uniqueness, if $Z_{n_{0}+n}$ satisfies (B.1), then $Z_{n_{0}+n}-Y_{n_{0}+n}=\beta^{2}(Z_{n_{0}+n-1}-Y_{n_{0}+n-1})$ for all $n>0$ with $Z_{n_{0}}-Y_{n_{0}}=0$. This implies $Z_{n_{0}+n}-Y_{n_{0}+n}=(Z_{n_{0}}-Y_{n_{0}})\,\beta^{2n}=0$ for all $n\geq 0$.
We can apply the previous claim and get $Z_{n_{0}+n}\leq-Y_{n_{0}+n}$. ∎
Proposition B.2.
Let $A\geq 0$, $X_{n}\geq 0$, $c_{n}\geq 0$.
Proof.
Define the sequence
$Y_{n}=A+\sum_{k=1}^{n}c_{k-1}\,X_{k-1}~~~,~~~Y_{0}=A$
$\frac{Y_{n}}{Y_{n-1}}\leq 1+c_{n-1}~~,~~\forall n\geq 1.$
Appendix C Some Facts Used in Proofs
Proposition C.1.
@@ Appendix C Some Facts Used in Proofs
Proof of (1). We use the same argument as (C.1), but replace $\ell(z,w_{*})$ with $F_{*}$.
Proof of (2). Use $F(w_{t})=F(w_{t})-F_{*}+F_{*}$ and apply lemma C.4.
Proof of (3.a). Applying lemma C.4 and subsequently Young’s inequality [8, eq 4.3] with $p=(1+\gamma)/(1-\gamma)$ and $q=(1+\gamma)/(2\gamma)$ for $\gamma\in(0,1]$ yields
$\|\nabla\ell(z,w)\|^{2}$ $\leq\left(\frac{1+\gamma}{\gamma}\right)^{\frac{2\gamma}{1+\gamma}}L^{\frac{2}{1+\gamma}}\left(\ell(z,w)-\ell(z,w_{*})\right)^{\frac{2\gamma}{1+\gamma}}$ (C.1)